OpenStudy (anonymous):

I solved a system of equations and the result was that it canceled out and i was left with just a number on the other side of the equal sign. Does that mean the answer is no solution or infinite solutions?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@zepdrix

OpenStudy (dan815):

can i see the 2 equations

OpenStudy (anonymous):

and if you get 0=0, does that mean no solution? or all real numbers (infinite solutions)

OpenStudy (mathmate):

Chances are, yes, if you end up with 4=9, or 0=5. Just to be sure, you may want to post the system of (how many) equations so that someone can help you check if it was a mistake somewhere.

OpenStudy (dan815):

yeah if the equalsign makes sense.. then infinite or else no solution

OpenStudy (anonymous):

alright! 3x+3y-7z=-2 4x+y-3z=8 -16x+5y-7z=-5

OpenStudy (anonymous):

those are the three equations, i get 36x-8z=104 -36x+8z=-45 at the end. and it cancels. you see what i mean?

OpenStudy (anonymous):

if it cancels to say, 4=5 then there is no solution but if it is like 0=0 it is always true and an identity meaning infinite

OpenStudy (anonymous):

well what if its like 0=56

OpenStudy (anonymous):

does 0 ever =56?

OpenStudy (anonymous):

noooooo

OpenStudy (anonymous):

exactly so it is never

OpenStudy (anonymous):

;) I'm just a tard. thank you

OpenStudy (aum):

36x-8z=104 -36x+8z=-45 Assuming the two equations you got are correct, 36x-8z=104 -36x+8z=-45 They can never be true for any value of x and y.

OpenStudy (mathmate):

Geometrically these two equations: 36x-8z=104 -36x+8z=-45 indicate two parallel lines with different y-intercepts. They are parallel and non-coinciding, so no solution.

OpenStudy (anonymous):

thank you guys!!!

OpenStudy (mathmate):

You're welcome! :)